# Linear response and exact hydrodynamic projections in Lindblad equations with decoupled Bogoliubov hierarchies

### v1

@@ VI.1.1 Eigenvalue spectrum
Figure 14: Real (red) and imaginary (blue) parts of the eigenvalues (106) as a function of total momentum for Model V with $\nu=1$, $\gamma=3$, $J=1$ and $L=100$. The green symbols show the purely real eigenvalues (105) for the same parameters.
@@
$J\gamma\cos\left(\frac{p}{2}\right)x^{3}-i\big{(}2J^{2}\cos\left(\frac{p}{2}\right)^{2}+\gamma^{2}\cos\left(p\right)\big{)}x^{2}+J\gamma\cos\left(\frac{p}{2}\right)\left(2\cos\left(p\right)+1\right)x+2iJ^{2}\cos\left(\frac{p}{2}\right)^{2}=0\ .$ (113)
@@
VII Operators cubic in fermions
@@
VIII Operators quartic in fermions
@@
Under time evolution these are expected to acquire hydrodynamic power-law tails at late times. In terms of the vectorized notations (13) the equations of motion for $\hat{\Psi}_{2,2}$ map onto a four-particle imaginary time Schrödinger equation. For Model I this can in principle still be solved exactly by using integrability methods [57]. However, in Models II, III and IV exact results can no longer be obtained following the same steps as in the two-particle sector, because the wave-functions of eigenstates of the Hamiltonians ${\cal H}$ no longer take simple forms in terms of superpositions of plane waves. This is because the two existing conserved quantities – momentum and the eigenvalue of ${\cal H}$ – are no longer sufficient to fix a set of four single-particle rapidities. As a result, in contrast to the situation in the two-particle sector, the wave functions of four-particle eigenstates in these models no longer takes (nested) Bethe ansatz form. We expect this to lead to quantifiably different behaviours of appropriate quantities related to $\hat{\Psi}_{2,2}(t)$ in Models II-IV compared to Model I. However, this line of enquiry goes beyond the scope of the present work and will be pursued elsewhere.
VIII.1 Model I: Bethe ansatz
As was noted in [21] the general eigenstates of ${\cal H}$ (15) in Model I have Bethe ansatz form and are parametrized in terms on $N=n_{\uparrow}+n_{\downarrow}$ rapidity variables $\{k_{j}|j=1,\dots,N\}$ and $M=n_{\downarrow}$ rapidity variables $\{\Lambda_{\alpha}|\alpha=1,\dots,M\}$. The two sets of rapidities fulfil the nested Bethe ansatz equations
@@ VIII.1 Model I: Bethe ansatz
The eigenvalue of ${\cal H}$ for a given solution is
@@
    $Lf_{1}(\lambda^{(1)}_{\alpha})$ $=2\pi J^{(1)}_{\alpha}+\sum_{\beta\neq\alpha}\theta\Big{(}\frac{\lambda^{(1)}_{\alpha}-\lambda^{(1)}_{\beta})}{\gamma}\Big{)}\ ,$ (126)
@@
In Fig. 15 we show exact diagonalization results for the eigenvalues of ${\cal H}$ in the 4-particle sector with $n_{\uparrow}=n_{\downarrow}=2$ in Model I, for parameter values $\gamma=3$, $J=1$ and $L=26$.
Figure 15: Real parts of the eigenvalues of ${\cal H}$ with $n_{\uparrow}=n_{\downarrow}=2$ as a function of total momentum for Model I with $\gamma=3$, $J=1$ and $L=26$. The black line is the eigenvalue of a length-2 $k-\Lambda$ string (class 1 in main text), and the gray area shows the continuum of two length-1 $k-\Lambda$ strings (class 2 in main text).
@@
In Fig. 16 we compare the eigenvalues obtained by using the string hypothesis as outlined above for $\gamma=2.5$, $J=1$ and $L=26$ to results obtained by exact diagonalization of the Hamiltonian ${\cal H}$ in the 4-particle sector.
Figure 16: Comparison of results obtained by the string hypothesis to exact diagonalization (circles) for the real parts of the eigenvalues as a function of total momentum for Model I with $\gamma=2.5$, $J=1$ and $L=26$. The four different classes of solutions (see main text) are shown as green, blue, and red dots.
We observe excellent agreement for all eigenvalues. This shows that it is possible to extend our analysis of section V to the four-particle sector and determine exact hydrodynamic projections of U(1) invariant operators quartic in fermions for Model I. However, this is beyond the scope of our present work.
VIII.2 Non-integrable models
In Models II-IV we still expect hydrodynamic modes, but as these models are not integrable a detailed analytic understanding of the eigenvalues of the various Hamiltonians ${\cal H}$ is not available. On the other hand, the numerical spectrum of eigenvalues can be straightforwardly obtained by exact diagonalization techniques.
Figure 17: Real parts of the eigenvalues of ${\cal H}$ in the sector $n_{\uparrow}=n_{\downarrow}=2$ as a function of total momentum for Model II with $\gamma=3$, $J=1$ and $L=26$. The gray area shows the continuum of two non-interacting $n_{\uparrow}=n_{\downarrow}=1$ ”excitations”.
As an example we show results for Model II with $\gamma=3$, $J=1$ and $L=26$ in Fig. 17. The eigenvalues with the largest real parts form a 2-particle scattering continuum, which we observe to approximately coincide with the continuum obtained by treating the eigenvalues in the $n_{\uparrow}=n_{\downarrow}=1$ sector (48) as corresponding to non-interacting ”excitations” and adding two of them in a restricted range of momenta $p$, which we fix by requiring that $\lambda(p)\geq-\gamma$. This requirement is motivated by the observation that other eigenvalues from a continuum of states starting at $-2\gamma$, cf. Fig. 17. It would be interesting to pursue a more precise description by treating the ”particles” as interacting and following the logic of Ref. [62].
VIII.3 Non $U(1)$ symmetric operators
@@
IX Non-equal time correlation functions in the steady state
@@
This has a simnple form in terms of wave-functions
@@
$G(q,\omega)=\sum_{q}\int_{0}^{\infty}\mathrm{d}t\ e^{i(\omega t-qx)}g(x,x|0,0;t)\ .$ (136)
@@
$G(q,\omega)\propto\frac{1}{i\omega+\alpha q^{2}}\ .$ (137)
Here the model dependent parameter $\alpha$ was previously obtained in our derivation of hydrodynamic projections in (87), (95), (96) and (97) respectively. In Figs 18 and 19 we show numerical results for ${\rm Im}\big{(}G(q,\omega)\big{)}$ in respectively Model I and II, for system size $L=500$ and parameters $J=1$ and $\gamma=3$.
@@
X Linear response in Lindblad dynamics
@@
[…]
+higher order in $\quad+\text{higher order in }\xi\ ,$ (141)
@@
X.1 Linear response in the steady state
The steady state of the unperturbed Limdblad evolution fulfils
@@ X.1 Linear response in the steady state
X.2 Linear response functions after a quantum quench
@@
×$\qquad\qquad\times\ \chi(j,\ell;t,t^{\prime})\ .$ (152)
@@
We stress that even in absence of dissipation, i.e. $\gamma=0$, the staggered dynamical suceptibilities $\chi^{0}_{\rm st}(Q,\omega,t^{\prime})$ vanish in the late time limit as a result of the restoration of translational invariance [49]. In Fig. 20 we show the real part of $\chi^{0}_{\mathrm{st}}(Q,\omega,t^{\prime}=0)$ as a function of $\omega$ and $Q$.
@@
In Fig. 21 we show the real part of $\chi^{0}_{\mathrm{st}}(Q,\omega,t^{\prime}=2)$ as a function of $\omega$ and $Q$. We observe that the magntiude of $\chi^{0}_{\mathrm{st}}$ diminishes with increasing $t^{\prime}$, in agreement with the fact that at late $t^{\prime}$ translational symmetry gets restored. We also observe an ”interference pattern” inside the particle-hole continuum.
@@ XI Summary and conclusions
Our work opens the door to several lines of enquiry. First, the decoupling of the BBGKY hierarchy holds on any lattice in any number of dimensions. It would be interesting to investigate whether explicit results e.g. for hydrodynamic projections can be obtained in higher dimensions. Second, we believe that the simpler structure of eigenstates of the aforementioned non-Hermitian Hamiltonians should lead to qualitative differences between integrable and non-integrable LEs in appropriated chosen complexity measures of time-evolved operators, cf. [64] for related work. This question is currently under investigation. Finally, it would be interesting to consider the effects of dissipative boundaries that preserve the truncation of the BBGKY hierarchy. This would enable the study of transport properties involving two-particle Green’s functions [65].
@@ Appendix A Diagonalizing the dual Lindbladian in momentum space
Here $\mathbf{e}_{\alpha}$ are unit vectors with a single entry at position $\alpha$. Finally, for Model IV we have introduced operators $T_{\alpha,\beta}$ that act on $\mathcal{O}_{n,m}$ as

### v2

@@ VII.1.1 Eigenvalue spectrum
Figure 14: Real (red) and imaginary (blue) parts of the eigenvalues (127) as a function of total momentum for Model V with $\gamma=3$, $J=1$ and $L=100$. The green symbols show the purely real eigenvalues (126) for the same parameters.
@@
$J\gamma\cos\left(\frac{p}{2}\right)x^{3}-i\big(2J^{2}\cos\left(\frac{p}{2}\right)^{2}+\gamma^{2}\cos\left(p\right)\big)x^{2}$
$+J\gamma\cos\left(\frac{p}{2}\right)\left(2\cos\left(p\right)+1\right)x+2iJ^{2}\cos\left(\frac{p}{2}\right)^{2}=0\ .$ (134)
@@
A physical process that could be analyzed in model V would be the time evolution of particle density and its fluctuations in a quantum quench starting from an initially empty state.
VIII Operators cubic in fermions
@@
IX Operators quartic in fermions
@@
[…]
IX.1 Model I: Bethe ansatz
[…]
@@ IX.1 Model I: Bethe ansatz
The eigenvalue ${\cal H} fo$r a given solution is
@@
    $Lf_{1}(\lambda^{(1)}_{\alpha})$ $=2\pi J^{(1)}_{\alpha}+\sum_{\beta\neq\alpha}\theta\Big(\frac{\lambda^{(1)}_{\alpha}-\lambda^{(1)}_{\beta}}{\gamma}\Big)\ ,$ (147)
@@
[…]
@@
[…]
Figure 16: Comparison of results obtained by the string hypothesis to exact diagonalization (circles) for the real parts of the eigenvalues as a function of total momentum for Model I with $\gamma=2.5$, $J=1$ and $L=26$. The four different classes of solutions (see main text) are shown as green, blue, red, and pink dots.
We observe excellent agreement for all eigenvalues. This shows that it is possible to extend our analysis of Section VI to the four-particle sector and determine exact hydrodynamic projections of U(1) invariant operators quartic in fermions for Model I. However, this is beyond the scope of our present work.
IX.2 Non-integrable models
[…]
As an example we show results for Model II with $\gamma=3$, $J=1$ and $L=26$ in Fig. 17. The eigenvalues with the largest real parts form a 2-particle scattering continuum, which we observe to approximately coincide with the continuum obtained by treating the eigenvalues in the $N_{\uparrow}=N_{\downarrow}=1$ sector (69) as corresponding to non-interacting ”excitations” and adding two of them in a restricted range of momenta $p$, which we fix by requiring that $\lambda(p)\geq-\gamma$. This requirement is motivated by the observation that other eigenvalues from a continuum of states starting at $2\gamma$, cf. Fig. 17. It would be interesting to pursue a more precise description by treating the ”particles” as interacting and extracting their ”scattering length” from the finite-size spectrum of the Lindbladian. The analogous analysis for the two-particle scattering continuum on the non-integrable spin-1 chain was carried out in Ref. [47].
IX.3 Non $U(1)$ symmetric operators
@@
X Non-equal time correlation functions in the steady state
@@
This has a simple form in terms of wave-functions
@@
$G(q,\omega)=\sum_{x}\int_{0}^{\infty}\mathrm{d}t\ e^{i(\omega t-qx)}g(x,x|0,0;t)\ .$ (157)
@@
$G(q,\omega)\propto\frac{1}{i\omega+Dq^{2}}\ .$ (158)
Here the model dependent parameter $D$ was previously obtained in our derivation of hydrodynamic projections in (108), (116), (117) and (118) respectively. In Figs 18 and 19 we show numerical results for ${\rm Im}\big(G(q,\omega)\big)$ in respectively Model I and II, for system size $L=500$ and parameters $J=1$ and $\gamma=3$.
@@
XI Linear response in Lindblad dynamics
@@
${\rm Tr}\left[A\rho(t)\right]-{\rm Tr}\left[A\rho_{0}(t)\right]$ $=\int_{0}^{\infty}dt^{\prime}\xi(t^{\prime})\chi_{A{\cal L}_{1}}(t,t^{\prime})$
$\quad+\text{higher order in }\xi\ ,$ (162)
@@
XI.1 Linear response in the steady state
The steady state of the unperturbed Lindblad evolution fulfils
@@ XI.1 Linear response in the steady state
XI.2 Linear response functions after a quantum quench
@@
$\qquad\qquad\times\ \chi(j,\ell;t,t^{\prime})\ .$ (173)
@@
We stress that even in absence of dissipation, i.e. $\gamma=0$, the staggered dynamical susceptibilities $\chi^{0}_{\rm st}(Q,\omega,t^{\prime})$ vanish in the late time limit as a result of the restoration of translational invariance [30]. In Fig. 20 we show the real part of $\chi^{0}_{\mathrm{st}}(Q,\omega,t^{\prime}=0)$ as a function of $\omega$ and $Q$.
@@
In Fig. 21 we show the real part of $\chi^{0}_{\mathrm{st}}(Q,\omega,t^{\prime}=2)$ as a function of $\omega$ and $Q$. We observe that the magnitude of $\chi^{0}_{\mathrm{st}}$ diminishes with increasing $t^{\prime}$, in agreement with the fact that at late $t^{\prime}$ translational symmetry gets restored. We also observe an ”interference pattern” inside the particle-hole continuum.
@@ XII Summary and conclusions
Our work opens the door to several lines of enquiry. First, the decoupling of the BBGKY hierarchy holds on any lattice in any number of dimensions. It would be interesting to investigate whether explicit results e.g. for hydrodynamic projections can be obtained in higher dimensions. Second, we believe that the simpler structure of eigenstates of the aforementioned non-Hermitian Hamiltonians should lead to qualitative differences between integrable and non-integrable LEs in appropriately chosen complexity measures of time-evolved operators, cf. [23] for related work. This question is currently under investigation. Finally, it would be interesting to consider the effects of dissipative boundaries that preserve the truncation of the BBGKY hierarchy [72, 73, 66, 40, 39, 31]. This would enable the study of transport properties involving two-particle Green’s functions.
@@ Appendix A Diagonalizing the dual Lindbladian in momentum space
Here $\mathbf{e}_{\alpha}$ are unit vectors with a single entry at position $\alpha$. Finally, for Model V we have introduced operators $T_{\alpha,\beta}$ that act on $\mathcal{O}_{n,m}$ as